OpenStudy (anonymous):

Can anyone help me with a math problem PLEASE? The math group is too crowded. Here's the problem: -12y - 10 = - 6y + 14

OpenStudy (anonymous):

-12y - 10 = -6y + 14 -6y - 10 = 14 (add 6 to -12) -6y = 24 (add 10 to 14) y = -4 (divide each side by -6)

OpenStudy (anonymous):

wrong group please post in the mathematics group.